OpenStudy (anonymous):

A point Q on a segment with endpoints A (2, −1) and C (4, 2) partitions the segment in a 3:1 ratio. Find Q. You must show all work to receive credit. i dont want the answer i want help solving it and explanation

OpenStudy (anonymous):

3:1 ratio is 4 pieces. We can find the length of each partition in the x and y directions. For x the partition length is \(\frac{ 4-2 }{ 4 }\) and for y it's \(\frac{ 2-(-1) }{ 4 }\). Start by simplifying those

OpenStudy (anonymous):

so x being 2/4 and y 3/4 ?

OpenStudy (anonymous):

yes. and you can reduce x to ½. Since Q is closer to C, you can subtract those from the coordinates of C. \[(4-\frac{ 1 }{ 2},2-\frac{ 3 }{ 4 })\]

OpenStudy (anonymous):

Alternatively, you could add them to A 3 times since A is 3 partitions away. Either way you'll get the same result.

OpenStudy (anonymous):

wait so am i solving whats what you just wrote ?

OpenStudy (anonymous):

2,1.35 ?

OpenStudy (anonymous):

(3.5,1.25) either you subtracted wrong or made a typo

OpenStudy (anonymous):

yes typo

OpenStudy (anonymous):

everything clear?

OpenStudy (anonymous):

so then i subtract ?

OpenStudy (anonymous):

no that's your answer.

OpenStudy (anonymous):

oh wow easier then i thought thank you so much !

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

want to help with another ? has to do with a graph and area